### v2

@@ 7   Conclusion
Instrumental variable strategies are pervasive in economics to estimate the effect of a treatment on an outcome. In this design, we commonly face two challenges: the internal and external validity of the IV estimate. To overcome at least the external validity, many studies leverage variation occurring from the different timing of policy adoption across units as an instrument for the treatment, and use instrumental variable techniques. The instrument, however, generally fails to satisfy the independence assumption in practice, and the underlying identification strategy is parallel those in DID designs.
@@
In this paper, we provide a new econometric framework, which we call an instrumented difference-in-differences (DID-IV). We start by considering a simple setting with two periods and two groups. In this setting, our DID-IV design mainly consists of a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between the two groups. We show that under $2\times 2$ DID-IV designs, the Wald-DID estimand captures the local average treatment effect on the treated (LATET). We also clarify the interpretation of the parallel trends assumption in the outcome. Specifically, we show that in DID-IV designs, time also plays the role of instrument, and the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups. Finally, we compare DID-IV to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and point out the issues inherent in Fuzzy DID designs.
@@
Next, we consider the DID-IV design in the case where there are more than two time periods and units become exposed to the instrument at different times. We call this a staggered DID-IV design, and formalize the target parameter and identifying assumptions. Specifically, our target parameter in staggered DID-IV designs is the cohort specific local average treatment effects on the treated (CLATT). The identifying assumptions are the natural generalization of those in $2\times 2$ DID-IV designs.
@@
We also provide the estimation method in staggered DID-IV designs, which does not require strong restrictions on treatment effect heterogeneity. In practice, most DID-IV applications implement the staggered DID-IV design by running TWFEIV regressions. In companion paper (Miyaji (2024)), however, we show that the conventional approach is inadequate for summarizing the treatment effects if the effect of the instrument on the treatment or the outcome is not stable over time. Our estimation method carefully chooses the comparison groups and does not suffer from the bias arising from the time-varying exposed effects. We also propose the weighting scheme in staggered DID-IV designs, and explain how one can conduct pretrends test to assess the validity of the parallel trends assumption in the treatment and the outcome.
@@
Finally, we illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006) who estimate returns to schooling in the United Kingdom, exploiting the timing variation of the introduction of school reforms between British and North Ireland. First, we assess the underlying DID-IV identification strategy in the author’s setting and check whether the identifying assumptions are plausible. Next, we estimate TWFEIV regressions and our stacked two stage least squares regressions in each dataset. We find that the TWFEV estimate turns out to be strictly negative, whereas our STS estimates are all positive. Finally, we estimate the summary causal parameter by exploiting our STS estimates. The estimate for the summary measure is also positive and indicates that returns to schooling are substantial in the U.K.
@@
Overall, this paper provides a new econometric framework for estimating the causal effects when the treatment adoption is potentially endogenous over time, but researchers can use variation in the timing of policy adoption across units as an instrument for the treatment. To avoid the issue of using TWFEIV estimators in this design, we also provide a reliable estimation method that is free from strong restrictions on treatment effect heterogeneity. Further developing alternative estimation methods and diagnostic tools will be a promising area for future research, facilitating the credibility of DID-IV research design in practice.
@@ Appendix A Proofs in section 2
Next, to show that the Wald-DID estimand equals to the LATET, it suffices to show that
@@
The first equality follows from Assumption 1 (No carryover assumption) and Assumption 2 (Exclusion restriction). The second equality follows from the simple manipulation. The third equality follows from Assumption 4. The forth equality follows from Assumption 6 (Parallel Trends Assumption in the outcome) and the simple calculation. The fifth equality follows from the Law of Iterated Expectations and Assumption 3.
@@
Here the first equality follows from the Law of Iterated Expectation and the second equality follows from the definition of $w_{g^{t},e}$ and $\Delta_{g^{t},e}$. Clearly, the weight $w_{g^{t},e}$ sum to one by construction.
Completing the proof.
@@ B.1   Non-binary, ordered treatment
For the numerator in the Wald DID estimand, we have:
@@
Here the first equality follows from the simple manipulation and Assumption 4. The second equality follows from Assumption 6.
@@
where the second equality follows from the fact: $\lambda_{1,0}^{1}-\lambda_{0,0}^{1}=1$.
We note that $\lambda_{1,j}^{1}\geq\lambda_{0,j}^{1}$ from Assumption 3 and that $\lambda_{1,j}^{1}$ and $\lambda_{0,j}^{1}$ can take only two values, zero or one. Therefore, we obtain:
@@
For the denominator in the Wald-DID estimand, by the similar argument, we first obtain:
@@
Here the first equality follows from the simple algebra and Assumption 4. The second equality follows from Assumption 5.
@@ B.2   Repeated cross sections
In this section, we assume repeated cross section settings and present our identification results.
@@ B.2.1 Two time periods
In two periods and two groups settings, our target parameter is the local average treatment effect in period $1$:
@@
Theorem 5 below shows that the Wald-DID estimand is equal to the local average treatment effect on the treated (LATET) in period $1$ under Assumptions 14-18. Here, we assume the binary treatment, but non-binary, ordered treatment case is easy to extend as in Theorem 3 (thus we omit it for brevity). In the proof, we omit the index $i$ to ease the notation.
@@
The second equality follows from the simple manipulation. The third equality follows from Assumption 16 and the forth equality follows from Assumption 15 and Assumption 17.
@@
The second equality follows from the simple manipulation. The third equality follows from Assumption 16. The forth equality follows from Assumption 18 and the simple calculation. The fifth equality follows from the Law of Iterated Expectations and Assumption 15.
@@ B.2.2 Multiple time periods
Let $T_{i}\in\{1,\dots,T\}$ denote the time period when unit $i$ is observed and let $Z_{i,t}$ denote the instrument status under $T_{i}=t$. We assume Assumption 7 in section 3, that is, we assume the staggered assignment of the instrument across units. In addition, we assume Assumption 14 (Exclusion restriction), which allows us to write $Y_{i}(D_{i},z)=Y_{i}(D_{i})$. Let $E_{i}\in\{1,\dots,T,\infty\}$ denote the cohort which unit $i$ belongs to. Let $Y_{i}(D_{i}^{e})$ and $D_{i}^{e}$ denote potential outcomes and potential treatment choices when unit $i$ belongs to cohort $e$. Let $D_{i,t}^{e}$ denote potential treatment choices under $T_{i}=t$ when unit $i$ belongs to cohort $e$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are independent and identically distributed (i.i.d).
In multiple time periods and multiple cohorts settings, our target parameter is the cohort specific local average treatment effect on the treated in period $T=e+l$:
@@
When repeated cross section data are available and we have multiple time periods, our staggered DID-IV designs consist of Assumptions 7, 10, 11 (in section 3), Assumption 14 (Exclusion restriction), and the following parallel trends assumptions.
@@ B.3   Introducing the treatment path in $2\times 2$ DID-IV designs
In this section, we discuss the DID-IV identification strategy in two periods with a binary treatment, introducing the treatment path in potential outcomes $Y_{i,t}(d,z)$. We first consider the two types of the no anticipation assumption on potential outcomes $Y_{i,t}(d,z)$ instead of imposing Assumption 1 (No carryover assumption). We then define the expected time gain and formalize the parallel trends assumption in the outcome as in Assumption 6 in this setting. Finally, we show that a Wald-DID estimand identifies the LATET’ (see section 4) under the suitable assumptions similar to the ones in section 2.3. Henceforth, we consider the same setting as in section 2.1 except that we do not adopt Assumption 1.
@@
Assumption 21 (No anticipation on the outcome among the $CM^{Z}\land NT^{T}$ in an exposed group).
@@
This assumption requires that the potential outcome at period 00 is same as the baseline outcome $Y_{i,0}((0,0),z)$ given the instrument path $z$, for those who are assigned to the instrument $z=(0,1)$ and belongs to the type $CM^{Z}\land NT^{T}$. We note that this assumption corresponds to the no anticipation assumption usually made in the common DID set up as in Roth et al. (2023) (see Table 3 and Table 4 in section 2.5) [^8] Here, it is inconsequential whether the instrument path $z$ is included or not in potential outcomes because we have $d=z$ in the common DID set up..
Assumption 22 (No anticipation on the outcome among the $CM^{Z}\land DF^{T}$ in an exposed group).
@@
This assumption requires that the potential outcome at period 00 is same as the baseline outcome $Y_{i,0}((1,0),z)$ given the instrument path $z$, for those who are assigned to the instrument $z=(0,1)$ and belongs to the type $CM^{Z}\land DF^{T}$. Assumption 22 is inherent to the DID-IV identification strategy when we introduce the treatment path.
We can rationalize Assumption 21 and 22 by noting that both the two Assumptions and the no anticipation assumption proposed by the recent DID literature restrict the anticipatory behavior of the units who belong to the exposed group and the compliers $CM^{Z}$.
We now make the parallel trends assumption in the outcome that is suitable for the set up we are considering here. We first define the similar notation as in section 2.4. Let $\bar{\Delta}_{g^{t},e}$ denote the expectation of the time trends in each groups in group $E_{i}=e$:
@@
Let $\bar{\Delta}_{0}$ and $\bar{\Delta}_{1}$ denote the expected time gain in unexposed and exposed groups respectively:
@@
where the definition of the weight $w_{g^{t},e}$ is same as in section 2.4.
As we previously noted in section 2.4, the parallel trends assumption in the outcome can be interpreted as requiring the homogeneous expected time gain between exposed and unexposed groups. The following assumption indeed make a such restriction in the current set up.
@@
The expected time gain between exposed and unexposed groups is same, that is, we have $\bar{\Delta}_{0}=\bar{\Delta}_{1}$.
@@
Suppose Assumption 2-5 and Assumption 21-23 holds. Then, a Wald-DID estimand corresponds to $LATET^{\prime}$, that is, we have
@@
First, we consider the denominator of the Wald-DID estimand. From the similar argument in the proof of Theorem 1, one can show the following result:
@@
Here the second equality holds from the fact that we can partition the compliers $CM^{Z}$ in an exposed group into $(CM^{Z}\land DF^{T})$ and $(CM^{Z}\land NT^{T})$ (see Table 1).
@@
We next decompose $E[Y_{1}-Y_{0}|E=0]$ as follows.
@@
Here the first equality follows from Table 2 and the second equality follows from the definition of the expected time gain $\bar{\Delta}_{0}$.
@@
Appendix C Proof in section 4
@@ C.1   Proof of Theorem 4
First, we prove the consistency of our SLS estimator. Recall that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ take the following form:
@@
We assume that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are well defined. More precisely, we assume that cohort $e$ is non-empty, that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}=e\}>0$ and there exist some sets of the unexposed cohort $U\subseteq\{u:e-1<u\leq T\}$ that are non-empty, i.e., $\sum_{i=1}^{n}\mathbf{1}\{E_{i}\in U\}>0$. This assumption guarantees that each Wald-DID estimator
@@
First we derive the probability limit of $\widehat{CLATT}_{e,e+l}$. We note that by the Law of Large Numbers and Slutsky’s theorem, we have
@@
By the similar arguments, we have
@@
From the Slutsky’s theorem, we have
@@
Lastly, it remains to show that the right hand side of (20) is equal to $CLATT_{e,e+l}$.
@@
The third equality follows from Assumption 8, 9 and 11. The forth equality follows from Assumption 13 and the fifth equality follows from Assumption 10.
For the denominator, by the similar calculations,
@@
Combining the result (21) with (22), we obtain the desirable result:
@@
We use the following fact repeatedly in the derivation. This fact is also found in de Chaisemartin and D’Haultfœuille (2018).

### v4

@@ 7   Conclusion
Instrumental variable strategies are pervasive in economics to estimate the effect of a treatment on an outcome. In this design, however, we commonly face two challenges: the internal and external validity of the IV estimate. To overcome at least external validity, many studies have leveraged variation occurring from the different timing of policy adoption across units as an instrument for treatment. The instrument based on these policy changes, however, generally fails to satisfy the randomization assumption in IV designs in practice, casting doubt on the internal validity of the IV estimate. Rather, the underlying identification strategy is parallel to DID designs, under which we identify the effect of the policy shock on the outcome, leveraging the variation of the exposure to that policy shock across units and over time.
@@
In this paper, we formalize the underlying identification strategy as an instrumented difference-in-differences (DID-IV) by combining the IV techniques with DID designs. We start with a simple two periods and two groups setting. In this setting, our DID-IV design mainly consists of a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between the two groups. We show that the Wald-DID estimand captures the LATET under $2\times 2$ DID-IV designs. We also clarify the interpretation of the parallel trends assumption in the outcome. Specifically, we show that in DID-IV settings, time also plays the role of instrument, and the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups. Finally, we compare DID-IV to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and point out the issues inherent in Fuzzy DID designs.
@@
Next, we consider the DID-IV design in more than two time periods with units being exposed to the instrument at different times. We call this a staggered DID-IV design, and formalize the target parameter and identifying assumptions. Our target parameter in this design is the cohort specific local average treatment effects on the treated (CLATT). The identifying assumptions are the natural generalization of those in $2\times 2$ DID-IV designs.
@@
We also provide the estimation method in staggered DID-IV designs, which does not require strong restrictions on treatment effect heterogeneity. In practice, most DID-IV applications implement the staggered DID-IV design by running TWFEIV regressions. In companion paper (Miyaji (2024)), however, we show that the conventional approach is inadequate for summarizing the treatment effects if the effect of the instrument on the treatment or the outcome is not stable over time. Our estimation method carefully chooses the comparison groups and does not suffer from the bias arising from the time-varying exposed effects. We also propose the weighting scheme in staggered DID-IV designs, and explain how one can conduct pre-trend tests to assess the validity of the parallel trends assumption in the treatment and the outcome.
@@
Finally, we illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006) who estimate returns to schooling in the UK, exploiting the timing variation of the introduction of school reforms between British and North Ireland. First, we assess the underlying DID-IV identification strategy in the author’s setting and check whether the identifying assumptions are plausible. Next, we estimate a TWFEIV regression and our stacked two stage least squares regressions in each dataset. The TWFEV estimate turns out to be strictly negative, whereas our STS estimates are all positive. Finally, we estimate the summary causal parameter by exploiting our STS estimates. The estimate for the summary measure is also positive and indicates that returns to schooling are substantial in the UK.
@@
Overall, this paper provides a new econometric framework for estimating the causal effects when the treatment adoption is potentially endogenous over time, but researchers can use variation in the timing of policy adoption across units as an instrument for the treatment. To avoid the issue of using the TWFEIV estimator under this design, we also provide a reliable estimation method that is free from strong restrictions on treatment effect heterogeneity. Further developing alternative estimation methods and diagnostic tools will be a promising area for future research, facilitating the credibility of DID-IV research design in practice.
@@ Appendix A Proofs in Section 2
Next, to show that the Wald-DID estimand equals the LATET, it suffices to show that
@@
The first equality follows from Assumption 1 (No carryover assumption) and Assumption 2 (Exclusion restriction). The second equality follows from the simple manipulation. The third equality follows from Assumption 4. The fourth equality follows from Assumption 6 (Parallel Trends Assumption in the outcome) and simple calculation. The fifth equality follows from the Law of Iterated Expectations and Assumption 3.
@@
The first equality follows from the Law of Iterated Expectation and the second equality follows from the definition of $w_{g^{t},e}$ and $\Delta_{g^{t},e}$. Clearly, the weight $w_{g^{t},e}$ sums to one by construction.
This completes the proof.
@@ B.1   Non-binary, ordered treatment
For the numerator in the Wald DID estimand, we have
@@
The first equality follows from the simple manipulation and Assumption 4. The second equality follows from Assumption 6.
@@
where the second equality follows from $\lambda_{1,0}^{1}-\lambda_{0,0}^{1}=1$.
Note that $\lambda_{1,j}^{1}\geq\lambda_{0,j}^{1}$ from Assumption 3 and that $\lambda_{1,j}^{1}$ and $\lambda_{0,j}^{1}$ can take only two values, zero or one. Therefore, we obtain:
@@
For the denominator in the Wald-DID estimand, by similar argument, we obtain:
@@
The first equality follows from simple algebra and Assumption 4. The second equality follows from Assumption 5.
@@ B.2   Repeated cross sections
Here, we present our identification results under repeated cross section settings.
@@ B.2.1 Two time periods
In a two-period and two-group setting, our target parameter is the local average treatment effect on the treated in period $1$:
@@
Theorem 5 below shows that the Wald-DID estimand is equal to the LATET in period $1$ under Assumptions 14-18. Here, we assume the binary treatment, but the non-binary, ordered treatment case is easy to extend as in Theorem 3 (thus, we omit it for brevity). In the proof, we omit the index $i$ to ease the notation.
@@
The second equality follows from the simple manipulation. The third equality follows from Assumption 16 and the fourth equality follows from Assumption 15 and Assumption 17.
@@
The second equality follows from the simple manipulation. The third equality follows from Assumption 16. The fourth equality follows from Assumption 18 and the simple calculation. The fifth equality follows from the Law of Iterated Expectations and Assumption 15.
@@ B.2.2 Multiple time periods
Let $T_{i}\in\{1,\dots,T\}$ denote the time period when unit $i$ is observed and let $Z_{i,t}$ denote the instrument status under $T_{i}=t$. We adopt Assumption 7 in Section 3; that is, we assume the staggered assignment of the instrument across units. In addition, we adopt Assumption 14 (Exclusion restriction), which allows us to write $Y_{i}(D_{i},z)=Y_{i}(D_{i})$. Let $E_{i}\in\{1,\dots,T,\infty\}$ denote the cohort which unit $i$ belongs to. Let $Y_{i}(D_{i}^{e})$ and $D_{i}^{e}$ denote potential outcomes and potential treatment choices when unit $i$ belongs to cohort $e$. Let $D_{i,t}^{e}$ denote potential treatment choices under $T_{i}=t$ when unit $i$ belongs to cohort $e$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are independent and identically distributed (i.i.d).
In multiple time period and multiple cohort settings, our target parameter is the cohort specific local average treatment effect on the treated in period $T=e+l$:
@@
When repeated cross section data are available and we have multiple time periods, the staggered DID-IV designs consist of Assumptions 7, 10, 11 (in Section 3), Assumption 14 (Exclusion restriction), and the following parallel trends assumptions.
@@ B.3   Introducing the treatment path in $2\times 2$ DID-IV designs
We discuss the DID-IV identification strategy in two periods with a binary treatment, introducing the treatment path in potential outcomes $Y_{i,t}(d,z)$. We consider two types of the no anticipation assumption on potential outcomes $Y_{i,t}(d,z)$ instead of imposing Assumption 1 (No carryover assumption). We then define the expected time gain and formalize the parallel trends assumption in the outcome as in Assumption 6 in this setting. Finally, we show that a Wald-DID estimand identifies the LATET’ (see Section 4) under suitable assumptions similar to the ones in Subsection 2.3. We consider the same setting as in Subsection 2.1 except that we do not adopt Assumption 1.
@@
Assumption 21 (No anticipation of the outcome among the $CM^{Z}\land NT^{T}$ in an exposed group).
@@
This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((0,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land NT^{T}$. Note that this assumption corresponds to the no anticipation assumption usually made in the common DID set up as in Roth et al. (2023) (see Table 3 and Table 4 in Subsection 2.5) [^9] It is inconsequential whether the instrument path $z$ is included or not in potential outcomes because we have $d=z$ in the common DID set up..
Assumption 22 (No anticipation of the outcome among the $CM^{Z}\land DF^{T}$ in an exposed group).
@@
This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((1,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land DF^{T}$. Assumption 22 is inherent to the DID-IV identification strategy when we introduce the treatment path.
We can rationalize Assumption 21 and 22 by noting that both of them and the no anticipation assumption proposed by the recent DID literature restrict the anticipatory behavior of the units belonging to the exposed group and the compliers $CM^{Z}$.
We now make the parallel trends assumption in the outcome suitable for the set up considered here. We first define the similar notation as in Subsection 2.4. Let $\bar{\Delta}_{g^{t},e}$ denote the expectation of the time trends in each groups in group $E_{i}=e$:
@@
Let $\bar{\Delta}_{0}$ and $\bar{\Delta}_{1}$ denote the expected time gain in unexposed and exposed groups, respectively:
@@
where the definition of the weight $w_{g^{t},e}$ is same as in Subsection 2.4.
As noted in Subsection 2.4, the parallel trends assumption in the outcome can be interpreted as requiring the homogeneous expected time gain between exposed and unexposed groups. The following assumption makes such a restriction in the current set up.
@@
The expected time gain between exposed and unexposed groups is the same; that is, we have $\bar{\Delta}_{0}=\bar{\Delta}_{1}$.
@@
Suppose Assumption 2-5 and Assumption 21-23 holds. Then, a Wald-DID estimand corresponds to $LATET^{\prime}$; that is, we have
@@
First, we consider the denominator of the Wald-DID estimand. From the argument in the proof of Theorem 1, one can show the following result:
@@
The second equality holds from the fact that we can partition the compliers $CM^{Z}$ in an exposed group into $(CM^{Z}\land DF^{T})$ and $(CM^{Z}\land NT^{T})$ (see Table 1).
@@
Next, we decompose $E[Y_{1}-Y_{0}|E=0]$ as follows.
@@
The first equality follows from Table 2 and the second follows from the definition of the expected time gain $\bar{\Delta}_{0}$.
@@
Appendix C Proofs in Section 4
@@ C.1   Proof of theorem 4
First, we prove the consistency of our SLS estimator. $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ take the following form:
@@
We assume that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are well defined. More precisely, we assume that cohort $e$ is non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}=e\}>0$ and there exist some sets of the unexposed cohort $U\subseteq\{u:e-1<u\leq T\}$ that are non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}\in U\}>0$. This assumption guarantees that each Wald-DID estimator
@@
First we derive the probability limit of $\widehat{CLATT}_{e,e+l}$. By the Law of Large Numbers and Slutsky’s theorem, we have
@@
By similar argument, we have
@@
From Slutsky’s theorem, we have
@@
Lastly, we show that the right hand side of (20) is equal to $CLATT_{e,e+l}$.
@@
The third equality follows from Assumption 8, 9 and 11. The fourth equality follows from Assumption 13 and the fifth follows from Assumption 10.
For the denominator, by similar calculations,
@@
Combining (21) with (22), we obtain the desirable result:
@@
The following fact, also found in de Chaisemartin and D’Haultfœuille (2018), is repeatedly used in the derivation.